This is part of the WFS activity. So far I have completed the following tasks:
1) I fixed the MEDM screens up to a point where they can be used for locking. There are still some buttons which invoke non-existing screens and some blank fields. But the basic filter banks and input and output matrices are fixed.
2) I copied all the old filter banks into the new screens both in the WFS head and in the WFS Master, where the servo filters are located. The I and Q filter banks in the WFS heads have been switched on.
3) I <=> Q phase settings in the WFS head for each quadrant: We have assumed that the I and Q are orthogonal so D=90 for all cases. I set the R phase to minimise the signal in all the Q lines. So the signal is largely in the I phase. I used Sine Response feature in DTT while supplying an excitation signal to MC2_ASCPIT_EXC. At times I used the YAW instead of PIT if I did not get enough coherence. This was set manually by watching the Q phase signal and minimising that by adjusting the R angle. It was in general possible to get this correct to a deg. There are several old scripts to do this in the MC/WFS but they do not work since most of them are based on the ezlockin or ezcademod functions. I will try to fix the ezWFS1phase and ezWFS2phase scripts to automate this. Some channel names have to be changed in these.
4) I measured the transfer function between the mirror motions [(MC1, MC2, MC3) x (PIT, YAW)] and the sensor DoF [(WFS1, WFS2) x (PIT, YAW)]. The measurements are reported below. The plan is to invert this matrix and use it as the Out_Matrix.
I list here the various steps I took in making this measurement.
a) Set the DC offsets on the individual quadrants to zero using an old script (which I updated with the new channel names). The script is called McWFS_dc_offsets and is located in the $scripts$/MC/WFS directory. Note that before doing this the PSL shutter was closed. This script sets a basic EPICS parameter called AOFF for each channel. These are listed in cvs/cds/caltech/target/c1iool0 .
b) Then the PSL beam into the MC was steered to optimise coupling into MC (described in my earlier post today). This is because we use the input beam as a reference while setting up the WFS.
c) Unlock the MC and center the directly reflected beam from the MC on the WFS. We use the DC monitors on the C1IOO_WFS_QPD.adl screen to center the spot on the WFS head.
d) Then used the WFSoffsets script to set the offsets in the I and Q filter banks to zero. This script uses the ezcaservo to look at the OUT16 channels and zeroes them by setting an appropriate offset. I took care to switch off all slow filters in the I and Q filter banks before this operation was carried out . Only the 60Hz comb filter was on.
e) Opened the PSL shutter and relocked the MC
f) Then I measured the transfer co-efs by oscillating the optic (exciting a specific degree of freedom) and observing the response in the WFS sensor degrees of freedom. These are tabulated above.
I plan to use this matrix and prepare the Output matix and then close the WFS servo loops.